### v1

# Operatorial formulation of crimo-taxis phenomena in a street

Abstract
In this paper, the ladder operators typical of quantum mechanics are used to construct an operatorial version of a model dubbed crimo-taxis. In a classical framework, the crimo-taxis model is described by reaction-diffusion partial differential equations describing a population divided into three interacting subgroups (ordinary citizens, drug users/dealers, and law enforcement personnel). By means of a quantum-like approach, these interacting subgroups are modeled using annihilation, creation, and number fermionic operators. Their dynamics is governed by a self-adjoint and time-independent Hamiltonian operator, emboding the interactions among the groups. Furthermore, a recent variant of the standard Heisenberg approach, namely $(\mathcal{H},\rho)$–induced dynamics, is also taken into account with the goal of enriching the dynamics. The results of some numerical simulations in a one–dimensional setting are presented and discussed.
Keywords. Operatorial models; $(\mathcal{H},\rho)$–induced dynamics; Heisenberg dynamics; Fermionic operators; Crimo-taxis phenomena.
Mathematics Subject Classification (2010). 37M05 - 37N20 - 47L90
1 Introduction
The dynamics of interacting species or populations is a crucial topic in theoretical biology, ecology, medicine, chemistry and other fields [1, 2].
In this paper, we focus on a special problem connected to urban crime, where different subgroups of a population can interact. In particular, our aim is to model the interaction among different subgroups of a population, i.e., ordinary citizens, drug users/dealers, and law enforcement personnel; as the crime phenomena do not occur uniformly in space and time, the various subgroups, due to their local interaction and diffusion processes, can spread in space at different times.
Traditionally, the dynamics of spatially distributed groups is modeled using reaction-diffusion partial differential equations [1, 2], which can exhibit interesting behaviors like pattern formation [3]. In particular, an original model, named crimo-taxis and proposed by Epstein in [1], involves three subgroups living in a one–dimensional spatial region representing a street. The original model does not possess asymptotically stable equilibria and this prevents to study Turing instability [4]; consequently, in [3], two possible variants have been proposed such that Turing instability is allowed and some interesting patterns emerge. The two models investigated in [3] can be written in compact form as
$\frac{\partial\mathbf{U}}{\partial t}=\mathbf{R}_{i}(\mathbf{U})+\mathcal{D}\frac{\partial^{2}\mathbf{U}}{\partial x^{2}},\qquad i=1,2,$ (1)
with
$\mathbf{U}=\left(\begin{array}[]{c}u\\ v\\ w\end{array}\right),\qquad\mathcal{D}=\left(\begin{array}[]{ccc}D_{11}&0&0\\ -D_{21}&D_{22}&D_{23}\\ -D_{31}&-D_{32}&D_{33}\end{array}\right),$
and $\mathbf{R}_{i}(\mathbf{U})$ ($i=1,2$) being one of the following reaction terms, say
$\mathbf{R}_{1}(\mathbf{U})=\left(\begin{array}[]{c}ru\left(1-\frac{u}{\kappa_{1}}\right)-\beta uv\\ \beta uv-\gamma vw\\ -bw+\xi uvw\end{array}\right),\qquad\mathbf{R}_{2}(\mathbf{U})=\left(\begin{array}[]{c}ru\left(1-\frac{u}{\kappa_{1}}\right)-\beta uv\\ \beta uv-\gamma vw\\ -bw+\xi uv\left(1-\frac{w}{\kappa_{2}}\right)\end{array}\right),$
for the first and second variant, respectively, whereas $u(x,t)$, $v(x,t)$, and $w(x,t)$ denote the densities of ordinary citizens (susceptibles), drug users/dealers (infectives), and law enforcement personnel at street position $x$ at time $t$, respectively; $r$ represents the susceptible reproduction rate, $\beta$ the infection rate, $\gamma$ the police arrest rate, $b$ the police natural decay rate, $\xi$ the police growth rate, whereas the coefficients $D_{ij}$ are the self- and cross-diffusion terms ($i=j$ and $i\neq j$, respectively). All parameters involved in (1) are assumed to be constant and positive (see [1, 3] for details).
Both models admit asymptotically stable homogeneous coexistence equilibria, susceptible of losing their stability due to the self- and cross-diffusive terms; moreover, the patterns exhibited by these two variants show a different spatial distribution of the three subgroups [3].
In this paper, we propose an alternative approach using the mathematical framework of quantum mechanics [5, 6], specifically raising, lowering operators and number representation. In particular, we implement an operatorial model where the actors of the system are represented by annihilation, creation and number fermionic operators whose evolution is ruled by a time–independent Hermitian Hamiltonian $\mathcal{H}$ within the recently introduced approach of $(\mathcal{H},\rho)$–induced dynamics (see [7], and references therein).
The operatorial approach, initially developed in quantum mechanics, has been successfully applied to several macroscopic systems outside physics (see [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24], and references therein, where the main motivations suggesting the use of the tools originally developed in a quantum context to describe classical situations have been widely discussed).
In this framework, we consider a system made by the three subgroups of a population in a patchy environment, that compete both locally and nonlocally, each one subject to migration phenomena with different mobilities; the actors of the model are represented by fermionic operators in a Hilbert space. This choice is advantageous for two main reasons: technical, because the Hilbert space is finite-dimensional, ensuring bounded operators, practical, because the mean values of the observables over an initial condition can represent the local densities of subgroups in different cells.
It is worth of being remarked that operatorial models for spatially distributed competing populations, where migration effects are included, have been already proposed [25, 26, 27]. However, our model description includes also non-local competition mechanisms (as in [24]). Moreover, in addition to the standard Heisenberg view to dynamics, we use the extension known as $(\mathcal{H},\rho)$–induced dynamics [7, 23, 28]. The latter approach studies the evolution of the system, as ruled by the Hamiltonian, but with the action, at fixed instants, of some rules whose effect is that of modifying the values of some parameters entering the Hamiltonian (without modifying the structure of the Hamiltonian) as a consequence of the evolution of the system itself. In such a way, the model adjusts itself during the evolution; the rule can be seen as a way to describe the change of the attitudes of the actors during the evolution, avoiding the introduction of a time dependent Hamiltonian or the consideration of other elements into the system such as a reservoir [29], with associated analytical and computational complications. In some applications (see [23, 28, 30]), the effect of the rules was to allow the system to approach asymptotic equilibrium states. However, we observe that this behavior does not necessarily manifest itself in the spatially distributed model investigated in this paper, even if it introduces a kind of irreversibility in the evolution.
The paper is structured as follows. In Section 2, we fix the notation and provide an operatorial formulation of the crimo-taxis model. The dynamics is governed by a time-independent self-adjoint quadratic Hamiltonian operator, which accounts for both local and nonlocal competition and migration effects. Furthermore, the analytical solution is exhibited. In Section 3, the $(\mathcal{H},\rho)$–induced dynamics approach is briefly reviewed, and two suitable rules acting on inertia parameters only, or on inertia and diffusion parameters, providing also a social interpretation, are introduced. In Section 4, in a one–dimensional domain, two scenarios characterized by different initial spatial distributions of the three subgroups have been considered, and some numerical results, both in the case of the standard Heisenberg dynamics and using the $(\mathcal{H},\rho)$–induced dynamics approach, are presented. Finally, Section 5 contains our concluding remarks, as well as some future extensions of the operatorial model presented in this paper.
2 Operatorial framework
In this Section, we set the notation and construct the operatorial formulation of the crimo-taxis model, starting from a general description of its components. We introduce the fermionic operators $a_{j,\alpha}$ and $a_{j,\alpha}^{\dagger}$ ($j=1,2,3$, $\alpha=1,\ldots,N$) to represent the actors of the model (ordinary citizens, drug users/dealers and law enforcement personnel) in the generic cell $\alpha$ of a one–dimensional region $\Omega=[0,L]$ representing a street of length $L$. We note that the symbol $\dagger$ denote the adjoint operation, while $a_{j,\alpha}$ and $a^{\dagger}_{j,\alpha}$ denote the annihilation and creation operators, respectively. Furthermore, we need to introduce the number operators [8]
$\hat{n}_{j,\alpha}=a_{j,\alpha}^{\dagger}a_{j,\alpha},\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N.$ (2)
The choice (that could seem too much restrictive) of considering a one–dimensional spatial region is not just (or only) a trick to simplify the analysis; however, migratory phenomena are often observed along well defined directions. In such circumstances, a model where the three subgroups of population are assumed spatially distributed in a one–dimensional region can result appropriate.
According to the formalism of second quantization, the fermionic operators satisfy the Canonical Anticommutation Relations (CAR):
$\{a_{j,\alpha},a_{k,\beta}\}=0,\qquad\{a^{\dagger}_{j,\alpha},a^{\dagger}_{k,\beta}\}=0,\qquad\{a_{j,\alpha},a^{\dagger}_{k,\beta}\}=\delta_{j,k}\,\delta_{\alpha,\beta}\,\mathcal{I},$ (3)
where $j,k=1,2,3$, $\;\alpha,\beta=1,\ldots,N$, $\;\{A,B\}:=AB+BA$ denotes the anticommutator between the operators $A$ and $B$, $\mathcal{I}$ is the identity operator, and $\delta_{i,j}$ represents the Kronecker symbol.
The Hilbert space $\mathbb{H}$, where the fermionic model representing the system $\mathcal{S}$ lives, is constructed as the linear span of the orthonormal set of vectors
$\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}=\left(\prod_{\alpha=1}^{N}(a_{1,\alpha}^{\dagger})^{n_{1,\alpha}}\right)\left(\prod_{\alpha=1}^{N}(a_{2,\alpha}^{\dagger})^{n_{2,\alpha}}\right)\left(\prod_{\alpha=1}^{N}(a_{3,\alpha}^{\dagger})^{n_{3,\alpha}}\right)\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}},$ (4)
where
$\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\equiv\varphi_{n_{1,1},\ldots,n_{1,N},n_{2,1}\ldots,n_{2,N},n_{3,1}\ldots,n_{3,N}},$ (5)
$\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}\equiv\varphi_{0,0,\ldots,0},$
i.e., the vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ is obtained by acting on the vacuum $\varphi_{\mathbf{0},\mathbf{0},\mathbf{0}}$ with the powers of the operators $a_{j,\alpha}^{\dagger}$, whit $n_{j,\alpha}=0,1$. As a consequence, we get $\hbox{dim}(\mathbb{H})=2^{3N}$.
The vector $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ means that it is initially assigned a mean value equal to $n_{j,\alpha}$ to the $j$–th population in the cell $\alpha$ such that
$\widehat{n}_{j,\alpha}\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}=n_{j,\alpha}\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},$ (6)
i.e., $\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}$ is the eigenvector of the number operator $\widehat{n}_{j,\alpha}$ associated to the eigenvalue $n_{j,\alpha}$.
The mean values of number operators, say
$n_{j,\alpha}=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},\widehat{n}_{j,\alpha}\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N$ (7)
are interpreted, once assigned an initial condition, as the local density of the population $\mathcal{P}_{j}$ in the cell $\alpha$.
The above operators are used to construct the Hamiltonian $\mathcal{H}$ associated to the system $\mathcal{S}$, which is the generator of the evolutive dynamics. In details, we will assume that
$\mathcal{H}=\mathcal{H}_{0}+\mathcal{H}_{I}+\mathcal{H}_{D}+\mathcal{H}_{C},$ (8)
where
$\mathcal{H}_{0}=\sum_{j=1}^{3}\sum_{\alpha=1}^{N}\omega_{j,\alpha}a_{j,\alpha}^{\dagger}a_{j,\alpha},$ (9)
$\mathcal{H}_{I}=\sum_{\alpha=1}^{N}\lambda_{1,\alpha}(a_{1,\alpha}a_{2,\alpha}^{\dagger}+a_{2,\alpha}a_{1,\alpha}^{\dagger})+\lambda_{2,\alpha}(a_{2,\alpha}a_{3,\alpha}^{\dagger}+a_{3,\alpha}a_{2,\alpha}^{\dagger}),$
$\mathcal{H}_{D}=\sum_{j=1}^{3}\sum_{\alpha=1}^{N}\left(\mu_{j,\alpha}\sum_{\beta=1}^{N}p_{\alpha,\beta}(a_{j,\alpha}a_{j,\beta}^{\dagger}+a_{j,\beta}a_{j,\alpha}^{\dagger})\right),$
$\mathcal{H}_{C}=\sum_{\alpha,\beta=1}^{N}p_{\alpha,\beta}\left(\nu_{1,\alpha}(a_{1,\alpha}a_{2,\beta}^{\dagger}+a_{2,\beta}a_{1,\alpha}^{\dagger})+\nu_{2,\alpha}(a_{2,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{2,\alpha}^{\dagger})+\right.$
$\qquad\qquad\left.+\nu_{3,\alpha}(a_{1,\alpha}a_{3,\beta}^{\dagger}+a_{3,\beta}a_{1,\alpha}^{\dagger})\right),$
are the different contributions in $\mathcal{H}$, whose description will be given below. In the expressions (9), the parameters $\omega_{j,\alpha}$, $\lambda_{j,\alpha}$, $\mu_{j,\alpha}$, and $\nu_{j,\alpha}$ ($\alpha=1,\ldots,N$, $j=1,\ldots,3$) are positive constants; moreover, the coefficients $p_{\alpha,\beta}$ ($\alpha,\beta=1,\ldots,N$), symmetric with respect to their indices, are expressed by
$p_{\alpha,\beta}=\begin{cases}1\quad\mbox{if }\beta\;\hbox{denotes a cell in the Moore neighborhood}\;M_{\alpha}\;\hbox{of cell}\;\alpha,\\ 0\quad\hbox{elsewhere}.\end{cases}$
Therefore, the cells in the Moore neighborhood of the cell $\alpha\neq 1,N$ are labeled as $\alpha-1$ and $\alpha+1$, whereas the cells labeled with 1 and $N$ have only one neighbor (labeled $2$ and $N-1$, respectively, see Figure 1).
Figure 1: Moore neighborhood of the cell $\alpha$.
Here, some comments about the various contributions (9) in the Hamiltonian $\mathcal{H}$ are detailed:
  • $\mathcal{H}_{0}$ is the free part of $\mathcal{H}$, and $\omega_{i,\alpha}$ are parameters somehow related to the inertia of the operators associated to the agents of $\mathcal{S}$: in fact, they can be thought of as a measure of the tendency of each degree of freedom to stay constant in time [8].
  • $\mathcal{H}_{I}$ rules the local (i.e., in the same cell $\alpha$) competitive interactions between the three subgroups. In particular, the coefficients $\lambda_{1,\alpha}$ ($\lambda_{2,\alpha}$, respectively) give a measure of the strength of the interactions between $a_{1,\alpha}$ and $a_{2,\alpha}$ ($a_{2,\alpha}$ and $a_{3,\alpha}$, respectively). The contribution $a_{1,\alpha}a_{2,\alpha}^{\dagger}$ is interpreted as a competition term since a particle associated to $a_{1,\alpha}$ is destroyed and a particle associated to $a_{2,\alpha}$ is created; the adjoint term $a_{2,\alpha}a_{1,\alpha}^{\dagger}$ swaps the roles of the two modes.
  • $\mathcal{H}_{D}$ is responsible for the diffusion of the three subgroups in the region; $\mu_{j,\alpha}$ represents the mobility coefficient of the $j$th population, $\mathcal{P}_{j}$, in the cell $\alpha$. For example, the contribution $a_{1,\alpha}a_{1,\beta}^{\dagger}$ is such that for the actor associated to $a_{1,\alpha}$ a particle is destroyed and for the actor associated to $a_{1,\beta}$ a particle is created (in other words, the particle is annihilated in cell $\alpha$ and created in cell $\beta$, where $\beta\in M_{\alpha}$, mimicking the diffusion in a neighboring cell); the adjoint term $a_{1,\beta}a_{1,\alpha}^{\dagger}$ swaps the roles of the two modes.
  • $\mathcal{H}_{C}$ takes into account a nonlocal competitive interaction of the three subgroups, and $\nu_{j,\alpha}$ measures the strength of nonlocal interaction (absent if $\nu_{j,\alpha}=0$); actually, the competition between the three populations occurs in adjacent cells.
We note also that $\mathcal{H}_{I}$ and $\mathcal{H}_{C}$ describe the effects of competition and cross-diffusion typically found in classical reaction-diffusion models [1].
Thence, adopting the Heisenberg representation, the operators $a_{j,\alpha}$ evolve in time according to the following ordinary differential equations:
$\dot{a}_{j,\alpha}=\textrm{i}[\mathcal{H},a_{j,\alpha}],\qquad j=1,\ldots,3,$ (10)
where $[A,B]=AB-BA$ denotes the the commutator between the operators $A$ and $B$. Since the Hamiltonian operator $\mathcal{H}$ is quadratic, it follows that the ordinary differential equations (10) turn out to be linear:
[…]
$\left.+\sum_{\beta=1}^{N}p_{\alpha\beta}((\mu_{1,\alpha}+\mu_{1,\beta})a_{1,\beta}+\nu_{1,\alpha}a_{2,\beta}+\nu_{3,\alpha}a_{3,\beta})\right),$
[…]
$+\left.\sum_{\beta=1}^{N}p_{\alpha\beta}\left((\mu_{2,\alpha}+\mu_{2,\beta})a_{2,\beta}+\nu_{1,\beta}a_{1,\beta}+\nu_{2,\alpha}a_{3,\beta}\right)\right),$
[…]
$+\left.\sum_{\beta=1}^{N}p_{\alpha\beta}((\mu_{3,\alpha}+\mu_{3,\beta})a_{3,\beta}+\nu_{2,\beta}a_{2,\beta}+\nu_{3,\beta}a_{1,\beta})\right).$
In the following Remark, we underline a possible analogy between reaction–diffusion equations and operatorial models.
Remark 1.
Under a suitable substitution it is possible to rewrite the system (11) in the following form
$\dot{a}_{1,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{1,\alpha}a_{1,\alpha}+2\nu_{3}a_{3,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{2,\alpha}+\widetilde{\mu}_{1}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right.$ (12)
$\qquad\left.+\widetilde{\nu_{1}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{2,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{2,\alpha}a_{2,\alpha}+\widetilde{\lambda}_{1,\alpha}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{3,\alpha}+\widetilde{\mu}_{2}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1}}{\Delta x^{2}}\right.$
$\qquad\left.+\widetilde{\nu_{1}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}+\widetilde{\nu_{2}}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right),$
$\dot{a}_{3,\alpha}=\textrm{i}\left(-\widetilde{\omega}_{3,\alpha}a_{3,\alpha}+2\nu_{3}a_{1,\alpha}+\widetilde{\lambda}_{2,\alpha}a_{2,\alpha}+\widetilde{\mu}_{3}\frac{a_{3,\alpha-1}-2a_{3,\alpha}+a_{3,\alpha+1}}{\Delta x^{2}}\right.$
$\qquad\left.+\widetilde{\nu_{2}}\frac{a_{2,\alpha-1}-2a_{2,\alpha}+a_{2,\alpha+1})}{\Delta x^{2}}+\widetilde{\nu_{3}}\frac{a_{1,\alpha-1}-2a_{1,\alpha}+a_{1,\alpha+1}}{\Delta x^{2}}\right),$
where
$\widetilde{\omega}_{i}=\omega_{i}-4\mu_{i},$ $\qquad\widetilde{\lambda}_{i,\alpha}=\lambda_{\alpha}+2\nu_{i},$ (13)
$\widetilde{\mu}_{i}=2\mu_{i}\Delta x^{2},$ $\qquad\widetilde{\nu_{i}}=\nu_{i}\Delta x^{2}.$
Particular emphasis is placed on the last three terms of each equation in the system (12), which can be interpreted as the spatial discretization of the second-order spatial derivative (with step size $\Delta x$). These terms represent the self-diffusion and cross-diffusion, respectively. Therefore, an operatorial model of a lattice can be thought of as the spatial discretization of a system of partial differential equations of reaction-diffusion type.
2.1 Analytical solution
The unknowns in (11) are operators represented by $2^{3N}\times 2^{3N}$ matrices with complex entries; whereupon, at least in principle, we have to solve a system of $3N\cdot 2^{6N}$ scalar ordinary differential equations in the complex domain. Nevertheless, because of the linearity of equation (11), we are able to reduce drastically the computational complexity. In fact, introducing the formal column vector
$\mathbf{A}=\left[a_{1,1},\ldots,a_{1,N},a_{2,1},\ldots,a_{2,N},a_{3,1},\ldots,a_{3,N}\right]^{T},$
(where the symbol T stands for the transposition operator) and a suitable $3N\times 3N$ real matrix $\Gamma$ (whose entries, once we fix $N$, can be constructed from (11)), the evolution equations for the annihilation operators can be written in the compact form as
$\frac{d\mathbf{A}}{dt}=\textrm{i}\Gamma\,\mathbf{A}.$ (14)
Thence, the formal solution of the system (14) is expressed by
$\mathbf{A}(t)=\exp\left(\textrm{i}\,\Gamma t\right)\mathbf{A}_{0}=B(t)\mathbf{A}_{0}.$ (15)
Here, we denote by $n_{i,\alpha}^{0}$ the density of population $\mathcal{P}_{i}$ in the cell $\alpha$ at the initial time $t_{0}=0$, and by $\mathbf{n}^{0}$ the corresponding initial values, say
$\mathbf{n}^{0}=\left({n_{1}^{0}},\ldots,{n_{3N}^{0}}\right),$
where $n_{k}^{0}$ is the initial density of $j$th population at the cell $\alpha$, with $k=(j-1)N+\alpha$.
Denoting with $B_{j,k}$ the generic entry of the $3N\times 3N$ matrix $B(t)$, and
$n_{j,\alpha}(t)=\langle\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}},a^{\dagger}_{j,\alpha}(t)a_{j,\alpha}(t)\varphi_{\mathbf{n}_{1},\mathbf{n}_{2},\mathbf{n}_{3}}\rangle,\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N,$
it is easy to derive, by using the CAR (3), the formula giving the mean values of the number operators at time $t$ [24]
$n_{j,\alpha}(t)=\sum_{\beta=1}^{3N}\left|B_{(j-1)N+\alpha,\beta}(t)\right|^{2}n_{\beta}^{0},\qquad j=1,\ldots,3,\qquad\alpha=1,\ldots,N.$ (16)
The quadratic form of the Hamiltonian has an immediate consequence: the dynamic behavior that we can obtain in each cell is at most quasiperiodic. Furthermore, due to the relation
$\left[\mathcal{H},\sum_{\alpha=1}^{N}\left(\widehat{n}_{1,\alpha}+\widehat{n}_{2,\alpha}+\widehat{n}_{3,\alpha}\right)\right]=0,$
the Hamiltonian possesses a first integral, expressing the fact that the sum of the densities of the three populations all over the cells of the domain is constant in time.
More complex outcomes (not necessarily quasiperiodic) could be recovered including in the Hamiltonian terms more than quadratic [9, 31]. Unfortunately, in such case we would be forced to solve numerically a very huge number of differential equations. Another strategy, with the goal of enriching the dynamics without rendering the problem computationally hard, has been recently introduced in [7, 23, 28, 32, 33], where interesting dynamics evolution is obtained still retaining a quadratic and time-independent Hamiltonian.
3 ($\mathcal{H},\rho$)–induced dynamics
The aim of the $(\mathcal{H},\rho)$–induced dynamics approach is to modify the standard Heisenberg dynamics by introducing some rules able to account for some effects that cannot be embedded in the definition of the Hamiltonian, unless we do not assume explicitly a time dependent Hamiltonian or consider an open quantum system including a reservoir, with consequent increase of technical difficulties (see [7]).
The concept of rule is nothing more than a law that modifies periodically some of the values of the parameters involved in the Hamiltonian as a consequence of the evolution of the system, so that the model is somehow able to adjust itself during the time evolution; essentially, this mathematical trick allows the actors to modify their attitudes and the strength of interactions according to the evolution of their state.
Hereafter, we sketch briefly the steps of the procedure (see [7] and references therein for further details):
  1. 1.
    divide the time interval $[0,T]$ in $n$ subintervals of length $\tau$;
  2. 2.
    let $k=1$, and consider a Hamiltonian operator $\mathcal{H}^{(k)}$;
  3. 3.
    using Heisenberg view, compute, in the time interval $[(k-1)\tau,k\tau]$, the evolution of ladder operators, whereupon, choosing an initial condition for the mean values of the number operators, obtain the time evolution of the latter (our observables);
  4. 4.
    according to the values of the observables at time $k\tau$, or to their variations in the time interval $[(k-1)\tau,k\tau]$, modify some of the parameters involved in $\mathcal{H}^{(k)}$;
  5. 5.
    a new Hamiltonian operator $\mathcal{H}^{(k+1)}$, having the same functional form as $\mathcal{H}^{(k)}$, but (in general) with different values of (some of) the involved parameters, is obtained;
  6. 6.
    increment by 1 the value of $k$;
  7. 7.
    if $k<n$ go to step 3.
From a mathematical point of view, the rule is just a map from $\mathbb{R}^{p}$ into $\mathbb{R}^{p}$ acting on the space of the $p$ parameters involved in the Hamiltonian. The global evolution is governed by a sequence of similar Hamiltonian operators, and the parameters entering the model can be considered stepwise (in time) constant. The complete evolution in the interval $[0,T]$ is obtained by glueing the local evolutions in each subinterval.
By using the $(\mathcal{H},\rho)$–induced dynamics approach, it is implicitly assuming to have a time dependent Hamiltonian. However, the time dependence is, in our case, of a very special form: in each interval [$(k-1)\tau,k\tau[$ the Hamiltonian does not depend on time, but in $k\tau$ some changes may occur, according to how the system is evolving. A comparison of this approach with that related to an explicitly time dependent Hamiltonian is discussed in [7].
In the model investigated in this paper, after fixing the value of $\tau$ (i.e., the length of the subintervals), we update at instants $k\tau$ ($k$ integer) the parameters as specified below.
Let us define
$\delta_{j,\alpha}^{(k)}=n_{j,\alpha}(k\tau)-n_{j,\alpha}((k-1)\tau),\qquad j=1,2,3,\quad\alpha=1,\ldots,N,$ (17)
$\delta_{\alpha}^{(k)}=n_{2,\alpha}(k\tau)n_{1,\alpha}(k\tau)-n_{2,\alpha}((k-1)\tau)n_{1,\alpha}((k-1)\tau),$
where $\delta_{j,\alpha}^{(k)}$ denotes the variation of the mean values of the number operators associated to each subgroup, i.e., how the subgroup densities vary at each step, and $\delta_{\alpha}^{(k)}$ is the difference between the products of the mean values of the number operators, i.e., it represents the variation of the interaction between ordinary citizens and drug addicts.
Now, let us consider the following rules for the three families of inertia and diffusion parameters, respectively, that come into play in the Hamiltonian by updating them at fixed times $k\tau$ ($k=1,2,\ldots)$, say
$\omega_{1,\alpha}\,\rightarrow\,\omega_{1,\alpha}\left(1-\delta_{2,\alpha}^{(k)}+\delta_{3,\alpha}^{(k)}\right),$ (18a)
$\omega_{2,\alpha}\,\rightarrow\,\omega_{2,\alpha}\left(1+\delta_{1,\alpha}^{(k)}-\delta_{3,\alpha}^{(k)}\right),$ (18b)
$\omega_{3,\alpha}\,\rightarrow\,\omega_{3,\alpha}\left(1-(\delta_{3,\alpha}^{(k)}+\delta_{\alpha}^{(k)})\right),$ (18c)
and
$\mu_{1,\alpha}\,\rightarrow\,\mu_{1,\alpha}\left(1+\delta_{2,\alpha}^{(k)}-\delta_{3,\alpha}^{(k)}\right),$ (19a)
$\mu_{2,\alpha}\,\rightarrow\,\mu_{2,\alpha}\left(1-\delta_{1,\alpha}^{(k)}+\delta_{3,\alpha}^{(k)}\right),$ (19b)
$\mu_{3,\alpha}\,\rightarrow\,\mu_{3,\alpha}\left(1-\delta_{\alpha}^{(k)}\right).$ (19c)
Some comments about the rule are in order. The relation (18a) indicates that the inertia parameter of the population $\mathcal{P}_{1}$ in the cell $\alpha$ increases (or decreases) when the difference between the local densities of police and drug addicts in the subinterval of length $\tau$ increases (or decreases). This suggests that a population increasing its inertia parameter tends to lower its tendency to change. The rationale of this rule is that an increase of the local density of this difference in the cell $\alpha$ has the effect of inducing a lower tendency to change (the police local density is greater than the one of drug addicts, so the population would feel safe), whereas a decrease induces the population to be less conservative (the local density of drug addicts is greater than the police one, so the population would feel less safe).
The relation (18b) states that the inertia parameter of the population $\mathcal{P}_{2}$ in the cell $\alpha$ increases (or decreases) as the difference between the local densities of the ordinary citizens and police increases (or decreases). This rule takes into account that an increase of the local density of this difference in the cell $\alpha$ has the effect of inducing a lower tendency to change (the population local density is greater than police one, therefore drug addicts continue to commit crime), whereas a decrease induces the drug addicts to be less conservative (the police local density is greater than the one of ordinary citizens, so that the drug addicts will tend to move to other areas to commit crimes).
The rule for inertia parameters of law enforcement personnel, as expressed in (18c), incorporates the concept of social alarm introduced by Epstein [1]. The inertia parameters are influenced by the sum of local densities of interactions between the ordinary citizens and drug addicts, alongside the local density of police. This social alarm is crucial for illustrating scenarios where law enforcement intervention is necessary to restore order.
The relation (19a) indicates that the diffusion parameter of the subgroup $\mathcal{P}_{1}$ in the cell $\alpha$ increases (or decreases) when the difference between the drug addicts and police local densities increases (or decreases). The idea behind it is that citizens must diffuse to neighboring territories when there is a high crime rate (the difference is positive), but if the police are sufficient in number to maintain order (the difference is negative), citizens can remain in their safe place.
The relation (19b) shows that the diffusion parameter of the drug addicts $\mathcal{P}_{2}$ in the cell $\alpha$ increases (or decreases) as the difference between the police and citizen local densities increases (or decreases). Drug addicts are less likely to dissolve when citizens density is high and police one is low, as this encourages continued criminal activity. In fact, with a high police presence, drug addicts may relocate to seek opportunities for crime elsewhere.
Finally, the relation (19c) reveals that the diffusion parameter of police $\mathcal{P}_{3}$ in the cell $\alpha$ increases (or decreases) as interactions between the drug addicts and citizens decrease (or increase). This implies that police should spread to other cells when social alarm in the cell $\alpha$ is low.
4 Numerical results
In this Section, we present some numerical simulations for the model (11).
In the first part, we neglect spatial inhomogeneities, i.e., all the interactions occur just in a single cell, whereas, the second one concerns with the model investigated in a one–dimensional spatial domain made of $N=50$ cells. In both situations, we compare the numerical results obtained by the standard Heisenberg dynamics and those coming from the superposition of the Heisenberg dynamics with the rules above described.
4.1 Model in one cell
In the standard Heisenberg dynamics, we see in Figure 2 that the evolution exhibits a never ending oscillatory trend, as expected. On the contrary, when the rule (18) is used, the amplitude of the oscillations tends to decrease and the mean densities of the three subgroups tends to reach an asymptotic equilibrium (see Figure 2). Furthermore, from Figure 2, an interpretation can be that law enforcement personnel are able to keep better drug users/dealers under control. Note also that the time needed to get a stationary behavior after the transient behavior depends on the values of the parameters.
Figure 2: Time evolution of the mean values of the number operators of the three subgroups. Only one cell is considered. No rule (Figure 2), and rule (18) with $\tau=4$ (Figure 2). The initial conditions in both cases are: $n_{1}^{0}=0.8$, $n_{2}^{0}=0.08$, $n_{3}^{0}=0.04$. The inertia and interaction parameters are $\omega_{1}=0.5$, $\omega_{2}=0.3$, $\omega_{3}=0.2$, and $\lambda_{1}=0.04$, $\lambda_{2}=0.025$, respectively.
4.2 Spatial model
Let us consider a street represented by a one–dimensional region $\mathcal{C}$ divided in $N=50$ adjacent cells, and investigate two scenarios characterized by different spatial distributions of the subgroups. The two situations are described as follows:
first scenario:
there are two areas with a high concentration of ordinary citizens, two suburbs with a small concentration of drug users/dealers, and a very small and dislocated area, with respect to the center of the street, where law enforcement personnel are concentrated. This scenario can be represented by the following initial mean densities:
$n_{1,\alpha}^{0}=0.8\left(\exp(-(\alpha-18)^{2}/32)+\exp(-(\alpha-31)^{2}/19)\right),$ (20)
$n_{2,\alpha}^{0}=0.08\left(\exp(-(\alpha-6)^{2}/44)+\exp(-(\alpha-24)^{2}/26)\right),$
$n_{3,\alpha}^{0}=0.04\exp(-(\alpha-35)^{2}/15),\qquad\alpha\in\mathcal{C};$
second scenario:
there is an area with a high concentration of ordinary citizens, and two small suburban areas occupied by drug users/dealers and law enforcement personnel, respectively. This scenario is similar to the one considered in [1]. We choose the following initial mean densities:
$n_{1,\alpha}^{0}=0.8(\exp(-(\alpha-25)^{2}/50),$ (21)
$n_{2,\alpha}^{0}=0.08(\exp(-(\alpha-8)^{2}/42),\qquad\qquad\alpha\in\mathcal{C},$
$n_{3,\alpha}^{0}=0.04\exp(-(\alpha-33)^{2}/17).$
The initial densities corresponding to the three subgroups for both scenarios are shown in Figure 3.
Figure 3: Initial conditions in a one–dimensional street made of $N=50$ cells for the first (20) and second (21) scenario, respectively.
Now, let us fix the initial parameters entering the model. We choose the following two sets of inertia parameters
$\omega_{1,\alpha}=0.5+0.25\left(\exp(-(\alpha-18)^{2}/32)+\exp(-(\alpha-31)^{2}/19)\right),$ (22)
$\omega_{2,\alpha}=0.2+0.1\left(\exp(-(\alpha-6)^{2}/44)+\exp(-(\alpha-24)^{2}/26)\right),$
$\omega_{3,\alpha}=0.3+0.15\exp(-(\alpha-35)^{2}/15),$
and
$\omega_{1,\alpha}=0.5+0.25\exp\left(-(\alpha-25)^{2}/50)\right),$ (23)
$\omega_{2,\alpha}=0.2+0.1\exp\left(-(\alpha-8)^{2}/42)\right),$
$\omega_{3,\alpha}=0.3+0.15\exp\left(-(\alpha-33)^{2}/17)\right),$
corresponding to the first and second scenario, respectively. This choice is motivated by the fact that each agent can be thought to have an inertia similar to the initial spatial distribution of the corresponding subgroup.
Finally, we assign the interaction, diffusion and cross-diffusion parameters (as choosen in the model described in [1, 3]), say
$\begin{array}[]{lll}\lambda_{1,\alpha}=0.04,&\qquad\mu_{1,\alpha}=0.07(1+\xi_{1}),&\qquad\nu_{1,\alpha}=0.006,\\ \lambda_{2,\alpha}=0.025,&\qquad\mu_{2,\alpha}=0.05(1+\xi_{1}),&\qquad\nu_{2,\alpha}=0.006,\\ \phantom{g}&\qquad\mu_{3,\alpha}=0.035(1+\xi_{2}),&\qquad\nu_{3,\alpha}=0.003,\end{array}$ (24)
where $\xi_{1}\in[0,0.01]$ and $\xi_{2}\in[0,0.002]$ are small random perturbations. The constants $\xi_{1}$ and $\xi_{2}$ are introduced such that all agents may have different initial mobilities.
Some comments about the numerical simulations are in order. Figures 4 and 5, where first and second scenario, (20) and (21), respectively, are considered, show the time evolution of the mean densities of the three spatially distributed populations, considering both the classical Heisenberg dynamics (Subfigures $(a),(b),(c)$) and the $(\mathcal{H},\rho)$–induced dynamics approach, where the rules (18) and (19) (Subfigures $(d),(e),(f)$ and $(g),(h),(i)$, respectively) are used.
In both scenarios, when classical Heisenberg dynamics is used, we observe an oscillatory trend for all population subgroups, as expected, since no rule is enforced. The results show similarities with Epstein’s model [1], despite the presence of oscillations. In [1], drug users/dealers tend to move toward areas with a higher concentration of ordinary citizens, as these locations offer greater opportunities for committing crimes; law enforcement, in turn, intervenes in response to the presence of drug users/dealers, attempting to reduce their numbers and restore order. This continuous cycle of interaction among criminals, citizens, and law enforcement can generate oscillations in both crime distribution and police intervention areas, similar to the oscillatory behavior observed in Figures 4 and 5. Even in the absence of explicit rules, the feedback among these actors-criminals moving toward citizens, citizens distancing themselves from drug users/dealers, and police reacting can explain the observed dynamics. The phenomenon can be interpreted as a cyclic sequence of phases where crime increases, peaks, and is subsequently countered by police intervention, forming a self-sustaining, dynamically evolving system. It is worth of being remarked that, since the initial inertia parameters have been choosen proportional to the peaks of the initial mean densities of the three subgroups, the tendency of ordinary citizens to remain in their affluent neighborhood is higher with respect to the case of using fixed values for the parameters across the entire domain, as commonly found in the literature [24, 27, 34]. This leads to a significant difference when the classical Heisenberg dynamics is applied: since the inertia parameters assume higher values in areas of higher population densities, the three subgroups show a stronger tendency to remain in their current state at the density peaks. This results in greater resistance to change, leading to the emergence of small stationary pattern formations, although for very short periods of time.
By means of the $(\mathcal{H},\rho)$–induced dynamics approach, we modify the standard Heisenberg dynamics with the rules introduced in Section 3, and the time evolution of the three subgroups accordingly changes (see Figures 4 and 5). By using the rule (18) (see Subfigures $(d),(e),(f)$), involving only modifications of the inertia parameters, the system exhibits a dynamics where oscillations are more controlled, and some patterns remain stable for short periods of time; this behaviour was not possible before. This trend represents a significant first step toward approaching the patterns formation found in [3]. Thus, the obtained outcome can be interpreted as a formation of spatially stationary patterns but dynamically evolving over time, where the population distribution appears stably structured in space but subject to variations and oscillations. This dynamic complexity could represent an intermediate step allowing the formation of more sophisticated and coherent patterns, similar to Turing ones [4], which balance spatial instabilities with time fluctuations. In Subfigures $(g),(h),(i)$ (see Figures 4 and 5), where rules (18) and (19) modify both inertia and diffusion parameters, we note that a trend similar to the case where only the rule (18), acting on inertia parameters, occurs. Anyway, the key difference is that, in this situation, law enforcement personnel are not only attracted towards areas with high criminals concentration but also to areas where ordinary citizens density increases. This fact implies that law enforcement personnel respond not only to zones with high criminal activity but also to densely populated areas, for instance in situations involving public events, demonstrations, or strikes. In other words, their actions are no longer exclusively focused on crime prevention and suppression but also on managing public safety in contexts where a high density of people may occur. This broader and multifaceted approach may reflect a strategy of law enforcement personnel aimed to maintain order and security in high-visibility situations susceptible of rising social risk, thus contributing to a more complex and varied dynamics compared to the model based solely on criminal influence. Additionally, we observe the emergence of stripe-like structures, although some small oscillations persist, which cannot be entirely suppressed as they are intrinsic to the construction of the operatorial model through the quantum-like approach.
Figure 4: First scenario: time evolution and spatial distribution of the three subgroups densities. Ordinary citizens (4, 4 and 4), drug users/dealers (4, 4 and 4), and law enforcement personnel (4, 4 and 4). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (20), (22) and (24).
Figure 5: Second scenario: time evolution and spatial distribution of the three subgroups densities. Ordinary citizens (5, 5 and 5), drug users/dealers (5, 5 and 5), and law enforcement personnel (5, 5 and 5). No-rule (upper), rule (18) (center) and rules (18)–(19) (bottom). The values of initial mean densities and parameters are given in (21), (23) and (24).
5 Conclusions
In this paper, we constructed an operatorial model based on fermionic ladder operators to describe the dynamical behavior of three subgroups of a population competing each other and able to diffuse in a one–dimensional spatial region made by a finite number of adjacent cells; the three populations compete locally (in the same cell) and nonlocally (in the adjacent cells). The mean values of the number operators associated to the three agents in a cell are interpreted as a measure of their local density in the cell. The dynamics is ruled by a self–adjoint and time independent quadratic Hamiltonian operator; consequently, adopting the classical Heisenberg viewpoint, the dynamics is ruled by linear ordinary differential equations, and its evolution exhibits a never ending oscillatory trend. To enrich the dynamics, we used the $(\mathcal{H},\rho)$–induced dynamics approach, and suitable choices for the rules are given and motivated. The rules can modify both the inertia and diffusion parameters; however, the cases where the rules act on the inertia parameters only, and on both inertia and the diffusion parameters are investigated. Two scenarios, characterized by different initial spatial distributions of the three subgroups, are considered, and some numerical results, together with their social interpretations, have been presented. The dynamical outcomes show that this model is compatible with the coexistence of the three competing populations in the same environment. The model is susceptible of various extensions and generalizations. For instance, we may include in the Hamiltonian operator effects due to cooperative interactions [34] between the three subgroups of a population, as well as construct a spatial model in a two–dimensional setting, with the aim of investigating if some pattern formations may emerge. Some of these extensions are currently planned and will be the subject of forthcoming papers.
Acknowledgements
This work has been carried out with the patronage of “Gruppo Nazionale per la Fisica Matematica” (GNFM) of “Istituto Nazionale di Alta Matematica” (INdAM). M.G. acknowledges financial support from Finanziamento del Programma Operativo Nazionale (PON) “Ricerca e Innovazione” 2014-2020 a valere sull’Asse IV “Istruzione e ricerca per il recupero”- Azione IV - Dottorati e contratti di ricerca su tematiche dell’innovazione, CUP J45F21001750007. M.G acknowledges financial support from PRIN 2022 “Nonlinear phenomena in low dimensional structures: models, simulations and theoretical aspects”, grant n. 2022TMW2PY. C.F.M. acknowledges financial support from “National Centre for HPC, Big Data and Quantum Computing (HPC)”, Codice progetto CN00000013 – SPOKE 10, CUP D43C22001240001.